# Human Mobility Datasets Enriched With Contextual and Social Dimensions

### v1

@@ 1. Introduction
The widespread adoption of location-aware devices and location-based social media has led to the generation of vast amounts of tracking data. These data typically capture the time-stamped positions of moving entities, often enriched with contextual semantic information, e.g., visited points of interest (POIs), means of transportation, and weather conditions. However, despite the growing volume of such data produced by location service providers, the research community continues to face challenges in accessing publicly available semantically enriched human mobility datasets. This scarcity arises from a combination of proprietary ownership, commercial restrictions, and strict privacy regulations (e.g., the European GDPR). Additionally, each added semantic layer increases the risk of re-identification (Gomes24, ; Monreale23, ) and the complexity of data integration, making public release both difficult and sensitive. The trajectory datasets available in the community often suffer from limitations such as being outdated or lacking semantic layers (e.g., Microsoft’s Geolife (zheng2011geolife, ), Porto Taxi dataset (taxiporto2015, )), or rely on spatiotemporal sparse check-ins with limited contextual information (e.g., Foursquare NYC/Tokyo (yang2019revisiting, )). To address these gaps, some researchers have turned to fully synthetic simulators (AndreasPatterns24, ; GoceData24, ; KappSurvey24, ; matsim2016, ), which, however, come with significant limitations in realism and applicability.
@@
  • Two human mobility datasets, publicly released on Zenodo (zenodo_dataset2025, ), derived from real GPS trajectories voluntarily shared via OpenStreetMap (OSM) (osm, ), enriched with different semantic layers, including raw contextual information (weather conditions) and inferred attributes (stops, moves, POIs, and transportation means).
@@
  • RDF (Resource Description Framework) representations of the datasets, in addition to the tabular ones, enabling semantic reasoning and LLMs integration. Furthermore, this format aligns with the FAIR (Findable, Accessible, Interoperable, Reusable) principles (fair_principles, ).
  • A reproducible and transparent pipeline for dataset creation, accompanied by publicly available code (matdataset_github_2025, ), allowing practitioners to generate customized versions of the datasets.
@@ 2. Methodology and Datasets
The methodology to create the semantically enriched trajectories datasets is schematized in Figure 1. The base layer of the figure represents the underlying data sources, specifically: (1) real-world raw data collected from OpenStreetMap (osm, ) and Meteostat (meteostat, ); (2) the inferred mobility patterns (including stops, moves, visited POIs, and transportation means) extracted through a combination of algorithms and heuristics; (3) synthetic social media posts generated through a large language model (LLM) to simulate realistic content generated by users during their movements. The final output consists of datasets represented in both tabular and RDF formats.
@@ 2.1. Data collection
Raw GPS trajectory data. Raw GPS trajectory data are sourced from OpenStreetMap [^1] Data available under the Open Database License (ODbL) (https://opendatacommons.org/licenses/odbl/1.0/) by selecting geographic bounding boxes over Paris and New York City and using the JOSM tool (josm, ) along with the OSM API. Due to the very large number of GPS traces available in New York City, we first partition its bounding box into a grid, retrieve data from individual grid cells, and subsequently merge the data. At this stage, we also perform some basic preprocessing, mainly to reduce data size: we discard (1) duplicated trajectories, (2) trajectories for which no temporal information is available, and (3) trajectories that are not associated with a specific user-ID. In OSM, each user account can have multiple trajectories; accordingly, for each user, we chose to merge all their trajectories, hence yielding one trajectory per user. In the end, for the city of Paris we obtain 1185 trajectories spanning August 2007 - February 2025, while for New York City we obtain 153180 trajectories spanning February 2008 - April 2025. Spatial distributions of these trajectories are visualized in Figure 2, focusing on high-density central areas for clarity. Timestamp conversions from UTC to the respective city time zones (Paris and New York City) are also performed. The bounding box values and the scripts used for downloading and preprocessing the GPS trajectories are publicly available in our GitHub repository (matdataset_github_2025, ).
@@
POI data. We download POI data from OSM via the OSM API (code available in our repository (matdataset_github_2025, )). To capture a wide range of different POIs, we consider the following macro-categories: amenity, which encompasses a wide range of facilities and services, such as banks, schools, toilets, and hospitals; shop, which covers retail establishments, including supermarkets, bakeries, and clothing stores; tourism, which includes tourist attractions and related services, like museums, hotels, and viewpoints; historic, which denotes sites of historical significance, such as castles, ruins, and monuments; and leisure, which includes green spaces, recreation, sport facilities, and more. The bounding boxes used to download the POI data are the same as those used to download the trajectories. For the city of Paris, we download a total of 109079 POIs, while for New York City we download a total of 66577 POIs.
Weather data. We download historical weather data from Meteostat (meteostat, ) via their API (code in our GitHub repository (matdataset_github_2025, )). We select the weather stations of Charles De Gaulle Airport for Paris and the John F. Kennedy Airport for New York City, assuming they represent conditions for the entire urban areas. We also perform some preprocessing to prepare the weather data for semantic enrichment, i.e., information concerning temperature and inferred sky conditions (e.g., sunny, rainy). Finally, the weather data has a daily frequency and spans January 1st 1990 - June 3rd 2025.
@@ 2.3. Semantic Enrichment of Trajectories
Our semantic enrichment workflow comprises two main steps: (1) trajectory segmentation and (2) enrichment of trajectory segments with multiple semantic dimensions (as shown in Figure 1). The tool we use to implement the workflow is MAT-Builder (lettich2023semantic, ). In the following, as well as in our public repositories (matdataset_github_2025, ; zenodo_dataset2025, ), we provide all the necessary details to understand, reproduce, and freely customize the semantic enrichment workflow.
Trajectory Segmentation. This operation takes as input a dataset of preprocessed raw trajectories and partitions every trajectory into stop and move segments using the classic stop-and-move criterion (spaccapietra2008conceptual, ). A stop segment represents the event in which an individual is staying within a small area for at least some time. A move segment represents the transition from one stop to another. These segments represent the basic units for enriching the trajectories. Finally, we report that we set the minimum duration of a stop to 10 minutes, while the maximum spatial radius a stop can have is set to 0.2 km. This yields 969 stops (mean duration: $193.96\pm 833.72$ minutes) and 1479 moves (mean distance covered: $9.45\pm 12.18$ km) for Paris, and 967 stops (mean duration: $36.31\pm 110.87$ minutes) and 19474 moves (mean distance covered: $8.76\pm 6.90$ km) for New York City.
@@ 2.4. Synthetic social media data generation
More precisely, posts are generated based on characteristics derived from the nearest POIs to trajectory stop segment centroids, primarily using POIs’ names and categories as indicative context for plausible user activities. Moreover, to introduce diversity and ensure balanced data, we use four synthetic user metadata categories: gender (male, female, other), age (18–24, 25–34, 35–44, 45–54, 55–64, 65+), ethnicity (White, Black, Hispanic, Asian, Other), and social network (Twitter, Instagram, Facebook, Tripadvisor). The generation prompt explicitly instructs the model to behave as a creative social media post generator, and that it must generate posts aligned to the information detailed above. The prompt also instructs the model to avoid generic sentence starters, e.g., “I just visited” or “Just spent”, which were overrepresented in preliminary attempts. We also prompted the model to include hashtags. For the generation parameters, we limit the output to 64 new tokens, set do_sample to true to avoid deterministic outputs, and use a temperature value of $0.9$ to encourage diverse, yet coherent generations. Both the code and the prompt that generate social media data are publicly available on our GitHub repository (matdataset_github_2025, ).
@@ 2.5. Representing Enriched Trajectories
To ensure usability and foster reproducible research, our datasets are released in two complementary formats: tabular and RDF. The tabular representation stores the enriched trajectories in several Pandas dataframes. For further details about this format, we refer the reader to our GitHub (matdataset_github_2025, ) and Zenodo (zenodo_dataset2025, ) repositories.
To support more advanced querying, semantic reasoning, and exploratory analysis of semantically enriched trajectories, we also provide an RDF-based knowledge graph representation. In this format, entities such as users, trajectories, stops, moves, and POIs are explicitly modeled, along with the relationships among them. More details on the ontology we used to support the Knowledge Graph representation can be found in (lettich2023semantic, ) and in our repositories (matdataset_github_2025, ; zenodo_dataset2025, ). Representing enriched trajectories in RDF enables SPARQL-based semantic queries and reasoning through triplestore technologies. Indeed, this transformation converts raw GPS traces into structured, context-aware mobility data that captures user routines, activities, movement patterns, and relationships with both the environment and social media. Knowledge graphs offer a formal, machine-readable representation of spatial relationships, enabling both symbolic reasoning and embedding-based approaches. By explicitly linking spatial entities, they enhance the ability of large language models to perform targeted searches. Furthermore, this representation adheres to the FAIR principles (fair_principles, ).
@@ 3. Potential Research Impacts
The resource (matdataset_github_2025, ; zenodo_dataset2025, ) presented in this paper supports research in human behavior analysis, urban planning, and algorithm development. It is especially useful for model validation, knowledge graph construction, and foundation model design. Its open, customizable pipeline ensures long-term utility for the research community. Spanning temporal, spatial, semantic, and social dimensions, its heterogeneous structure captures the complexity of urban mobility. The semi-synthetic nature makes it well-suited for experimental tasks such as movement prediction, behavioral modeling, and semantic reasoning.
@@
While the classical tabular format is well-suited for machine learning pipelines, the RDF representation adds significant value by enabling knowledge graph construction and semantic reasoning (UrbanKG, ; MobilityPredictionKG, ). It offers a rich framework for linking structured mobility data with contextual semantic layers, supporting advanced and interpretable analyses like semantic enrichment, spatiotemporal reasoning, user-POI linking, and natural language processing for geographical data. This aspect could be useful in combination with LLMs to capture spatial semantics implicitly from large text corpora. Combining these approaches holds strong potential: knowledge graphs provide a reliable spatial structure, while LLMs bring powerful natural language understanding and generation capabilities. For instance, it can fuel emerging research on chatbot architectures capable of answering complex, natural language questions about mobility data (padoan2024mobility, ; chen2024travelagent, ; devunuri2024chatgpt, ). These systems often rely on multiple specialized agents – each driven by an LLM – that collaboratively query diverse data sources and process the retrieved information to formulate answers. Recent studies have already demonstrated the automatic generation of SPARQL queries over knowledge graphs (kovriguina2023sparqlgen, ), which could be directly applied to the semantic layers of our dataset. Another promising role for this resource is in the context of foundation models and LLMs. By publicly releasing these datasets and the building pipeline, we aim to empower the CIKM community with a resource that enables experiments and testing of methods and algorithms for more interpretable, explainable, and context-aware approaches to human behavior analysis.

### v2

@@ 1. Introduction
The widespread adoption of location-aware devices and location-based social media has led to the generation of vast amounts of tracking data. These data typically capture the time-stamped positions of moving entities, often enriched with contextual semantic information, e.g., visited points of interest (POIs), means of transportation, and weather conditions. However, despite the growing volume of such data produced by location service providers, the research community continues to face challenges in accessing publicly available semantically enriched human mobility datasets. This scarcity arises from a combination of proprietary ownership, commercial restrictions, and strict privacy regulations (e.g., the European GDPR). Additionally, each added semantic layer increases the risk of re-identification (Gomes24; Monreale23) and the complexity of data integration, making public release both difficult and sensitive. The trajectory datasets available in the community often suffer from limitations such as being outdated or lacking semantic layers (e.g., Microsoft’s Geolife (zheng2011geolife), Porto Taxi dataset (taxiporto2015)), or rely on spatiotemporal sparse check-ins with limited contextual information (e.g., Foursquare NYC/Tokyo (yang2019revisiting)). To address these gaps, some researchers have turned to fully synthetic simulators (AndreasPatterns24; GoceData24; KappSurvey24; matsim2016), which, however, come with significant limitations in realism and applicability.
@@
  • Two human mobility datasets, publicly released on Zenodo (zenodo_dataset2025), derived from real GPS trajectories voluntarily shared via OpenStreetMap (OSM) (osm), enriched with different semantic layers, including raw contextual information (weather conditions) and inferred attributes (stops, moves, POIs, and transportation means).
@@
  • RDF (Resource Description Framework) representations of the datasets, in addition to the tabular ones, enabling semantic reasoning and LLMs integration. Furthermore, this format aligns with the FAIR (Findable, Accessible, Interoperable, Reusable) principles (fair_principles).
  • A reproducible and transparent pipeline for dataset creation, accompanied by publicly available code (matdataset_github_2025), allowing practitioners to generate customized versions of the datasets.
@@ 2. Methodology and Datasets
The methodology to create the semantically enriched trajectories datasets is schematized in Figure 1. The base layer of the figure represents the underlying data sources, specifically: (1) real-world raw data collected from OpenStreetMap (osm) and Meteostat (meteostat); (2) the inferred mobility patterns (including stops, moves, visited POIs, and transportation means) extracted through a combination of algorithms and heuristics; (3) synthetic social media posts generated through a large language model (LLM) to simulate realistic content generated by users during their movements. The final output consists of datasets represented in both tabular and RDF formats.
@@ 2.1. Data collection
Raw GPS trajectory data. Raw GPS trajectory data are sourced from OpenStreetMap [^1] Data available under the Open Database License (ODbL) (https://opendatacommons.org/licenses/odbl/1.0/) by selecting geographic bounding boxes over Paris and New York City and using the JOSM tool (josm) along with the OSM API. Due to the very large number of GPS traces available in New York City, we first partition its bounding box into a grid, retrieve data from individual grid cells, and subsequently merge the data. At this stage, we also perform some basic preprocessing, mainly to reduce data size: we discard (1) duplicated trajectories, (2) trajectories for which no temporal information is available, and (3) trajectories that are not associated with a specific user-ID. In OSM, each user account can have multiple trajectories; accordingly, for each user, we chose to merge all their trajectories, hence yielding one trajectory per user. In the end, for the city of Paris we obtain 1185 trajectories spanning August 2007 - February 2025, while for New York City we obtain 153180 trajectories spanning February 2008 - April 2025. Spatial distributions of these trajectories are visualized in Figure 2, focusing on high-density central areas for clarity. Timestamp conversions from UTC to the respective city time zones (Paris and New York City) are also performed. The bounding box values and the scripts used for downloading and preprocessing the GPS trajectories are publicly available in our GitHub repository (matdataset_github_2025).
@@
POI data. We download POI data from OSM via the OSM API (code available in our repository (matdataset_github_2025)). To capture a wide range of different POIs, we consider the following macro-categories: amenity, which encompasses a wide range of facilities and services, such as banks, schools, toilets, and hospitals; shop, which covers retail establishments, including supermarkets, bakeries, and clothing stores; tourism, which includes tourist attractions and related services, like museums, hotels, and viewpoints; historic, which denotes sites of historical significance, such as castles, ruins, and monuments; and leisure, which includes green spaces, recreation, sport facilities, and more. The bounding boxes used to download the POI data are the same as those used to download the trajectories. For the city of Paris, we download a total of 109079 POIs, while for New York City we download a total of 66577 POIs.
Weather data. We download historical weather data from Meteostat (meteostat) via their API (code in our GitHub repository (matdataset_github_2025)). We select the weather stations of Charles De Gaulle Airport for Paris and the John F. Kennedy Airport for New York City, assuming they represent conditions for the entire urban areas. We also perform some preprocessing to prepare the weather data for semantic enrichment, i.e., information concerning temperature and inferred sky conditions (e.g., sunny, rainy). Finally, the weather data has a daily frequency and spans January 1st 1990 - June 3rd 2025.
@@ 2.3. Semantic Enrichment of Trajectories
Our semantic enrichment workflow comprises two main steps: (1) trajectory segmentation and (2) enrichment of trajectory segments with multiple semantic dimensions (as shown in Figure 1). The tool we use to implement the workflow is MAT-Builder (lettich2023semantic). In the following, as well as in our public repositories (matdataset_github_2025; zenodo_dataset2025), we provide all the necessary details to understand, reproduce, and freely customize the semantic enrichment workflow.
Trajectory Segmentation. This operation takes as input a dataset of preprocessed raw trajectories and partitions every trajectory into stop and move segments using the classic stop-and-move criterion (spaccapietra2008conceptual). A stop segment represents the event in which an individual is staying within a small area for at least some time. A move segment represents the transition from one stop to another. These segments represent the basic units for enriching the trajectories. Finally, we report that we set the minimum duration of a stop to 10 minutes, while the maximum spatial radius a stop can have is set to 0.2 km. This yields 969 stops (mean duration: $193.96\pm 833.72$ minutes) and 1479 moves (mean distance covered: $9.45\pm 12.18$ km) for Paris, and 967 stops (mean duration: $36.31\pm 110.87$ minutes) and 19474 moves (mean distance covered: $8.76\pm 6.90$ km) for New York City.
@@ 2.4. Synthetic social media data generation
More precisely, posts are generated based on characteristics derived from the nearest POIs to trajectory stop segment centroids, primarily using POIs’ names and categories as indicative context for plausible user activities. Moreover, to introduce diversity and ensure balanced data, we use four synthetic user metadata categories: gender (male, female, other), age (18–24, 25–34, 35–44, 45–54, 55–64, 65+), ethnicity (White, Black, Hispanic, Asian, Other), and social network (Twitter, Instagram, Facebook, Tripadvisor). The generation prompt explicitly instructs the model to behave as a creative social media post generator, and that it must generate posts aligned to the information detailed above. The prompt also instructs the model to avoid generic sentence starters, e.g., “I just visited” or “Just spent”, which were overrepresented in preliminary attempts. We also prompted the model to include hashtags. For the generation parameters, we limit the output to 64 new tokens, set do_sample to true to avoid deterministic outputs, and use a temperature value of $0.9$ to encourage diverse, yet coherent generations. Both the code and the prompt that generate social media data are publicly available on our GitHub repository (matdataset_github_2025).
@@ 2.5. Representing Enriched Trajectories
To ensure usability and foster reproducible research, our datasets are released in two complementary formats: tabular and RDF. The tabular representation stores the enriched trajectories in several Pandas dataframes. For further details about this format, we refer the reader to our GitHub (matdataset_github_2025) and Zenodo (zenodo_dataset2025) repositories.
To support more advanced querying, semantic reasoning, and exploratory analysis of semantically enriched trajectories, we also provide an RDF-based knowledge graph representation. In this format, entities such as users, trajectories, stops, moves, and POIs are explicitly modeled, along with the relationships among them. More details on the ontology we used to support the Knowledge Graph representation can be found in (lettich2023semantic) and in our repositories (matdataset_github_2025; zenodo_dataset2025). Representing enriched trajectories in RDF enables SPARQL-based semantic queries and reasoning through triplestore technologies. Indeed, this transformation converts raw GPS traces into structured, context-aware mobility data that captures user routines, activities, movement patterns, and relationships with both the environment and social media. Knowledge graphs offer a formal, machine-readable representation of spatial relationships, enabling both symbolic reasoning and embedding-based approaches. By explicitly linking spatial entities, they enhance the ability of large language models to perform targeted searches. Furthermore, this representation adheres to the FAIR principles (fair_principles).
@@ 3. Potential Research Impacts
The resource (matdataset_github_2025; zenodo_dataset2025) presented in this paper supports research in human behavior analysis, urban planning, and algorithm development. It is especially useful for model validation, knowledge graph construction, and foundation model design. Its open, customizable pipeline ensures long-term utility for the research community. Spanning temporal, spatial, semantic, and social dimensions, its heterogeneous structure captures the complexity of urban mobility. The semi-synthetic nature makes it well-suited for experimental tasks such as movement prediction, behavioral modeling, and semantic reasoning.
@@
While the classical tabular format is well-suited for machine learning pipelines, the RDF representation adds significant value by enabling knowledge graph construction and semantic reasoning (UrbanKG; MobilityPredictionKG). It offers a rich framework for linking structured mobility data with contextual semantic layers, supporting advanced and interpretable analyses like semantic enrichment, spatiotemporal reasoning, user-POI linking, and natural language processing for geographical data. This aspect could be useful in combination with LLMs to capture spatial semantics implicitly from large text corpora. Combining these approaches holds strong potential: knowledge graphs provide a reliable spatial structure, while LLMs bring powerful natural language understanding and generation capabilities. For instance, it can fuel emerging research on chatbot architectures capable of answering complex, natural language questions about mobility data (padoan2024mobility; chen2024travelagent; devunuri2024chatgpt). These systems often rely on multiple specialized agents – each driven by an LLM – that collaboratively query diverse data sources and process the retrieved information to formulate answers. Recent studies have already demonstrated the automatic generation of SPARQL queries over knowledge graphs (kovriguina2023sparqlgen), which could be directly applied to the semantic layers of our dataset. Another promising role for this resource is in the context of foundation models and LLMs. By publicly releasing these datasets and the building pipeline, we aim to empower the CIKM community with a resource that enables experiments and testing of methods and algorithms for more interpretable, explainable, and context-aware approaches to human behavior analysis.
Acknowledgments
This research has been partially funded by the European Union’s Horizon Europe research and innovation program EFRA (Grant Agreement Number 101093026). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or European Commission-EU. Neither the European Union nor the granting authority can be held responsible for them.